OpenStudy (anonymous):

1/(x-3)^2 - 2/(x-3) =15

OpenStudy (lalaly):

\[\large{\frac{1}{(x-3)^2}}-(\large{\frac{2}{x-3} \times \frac{x-3}{x-3}}) = \large{\frac{1-2x+6}{(x-3)^2}}\]

OpenStudy (lalaly):

\[=\large{\frac{7-2x}{(x-3)^2}}\]

OpenStudy (anonymous):

\[\frac{1-2(x-3)}{(x-3)^2}=15\] \[\frac{1-2x+6}{x^2+9-6x}=15\] \[7-2x=15(x^2+9-6x)\] \[7-2x=15x^2+135-90x\] \[7-2x+90x-135=15x^2\] \[88x-128=15x^2\] \[88x-15x^2=128\]

OpenStudy (lalaly):

\[\large{7-2x}=\large{15 \times (x-3)^2}\]

OpenStudy (anonymous):

\[\frac{8}{3},\frac{16}5\]